The length of a rectangular garden is 4 cm greater than the width. The area of the garden is 96 cm . Find the length and the width.
Answer by checkley77(12844)

L=W+4
LW=96
(W+4)W=96
W^2+4W-96=0
(W+12)(W-8)=0
W-8=0
W=8 ANS. FOR THE WIDTH.
L=8+4=12 ANS. FOR THE LENGTH.
PROOF:
8*12=96
96=96